OpenStudy (toxicsugar22):

A school of tuna is making a 200 mile trip at an average air speed of 3 miles per hour in still water, but is swimming againist a current. Express the tuna's travel time t, as a function of the current speed, v.

OpenStudy (anonymous):

So, their actual speed is the "air speed" minus the current: s = 3-v And the time it will take them to make the trip is the distance divided by their speed: t = 200/s t= 200/(3-v)

OpenStudy (toxicsugar22):

and write and solve an equation to find the current spees if the school make its trip in 8 hours.

OpenStudy (anonymous):

For that, just take the final equation and solve it for v: t = 200/(3-v) 3 - v = 200/t v = 3 - 200/t and plug in 8 for t: v = 3 - 200/8 = -22 (the negative number just means that the current is working in their favor)

OpenStudy (toxicsugar22):

so thats the answer

OpenStudy (toxicsugar22):

-22

OpenStudy (toxicsugar22):

is this right

OpenStudy (toxicsugar22):

can u help

OpenStudy (anonymous):

Should be right, assuming I didn't make a math error somewhere.